# OpenWorldSAM: Extending SAM2 for Universal Image Segmentation with Language Prompts

### v1

@@ Abstract
The ability to segment objects based on open-ended language prompts remains a critical challenge, requiring models to ground textual semantics into precise spatial masks while handling diverse and unseen categories. We present OpenWorldSAM, a framework that extends the prompt-driven Segment Anything Model v2 (SAM2) to open-vocabulary scenarios by integrating multi-modal embeddings extracted from a lightweight vision-language model (VLM). Our approach is guided by four key principles: i) Unified prompting: OpenWorldSAM supports a diverse range of prompts, including category-level and sentence-level language descriptions, providing a flexible interface for various segmentation tasks. ii) Efficiency: By freezing the pre-trained components of SAM2 and the VLM, we train only 4.5 million parameters on the COCO-stuff dataset, achieving remarkable resource efficiency. iii) Instance Awareness: We enhance the model’s spatial understanding through positional tie-breaker embeddings and cross-attention layers, enabling effective segmentation of multiple instances. iv) Generalization: OpenWorldSAM exhibits strong zero-shot capabilities, generalizing well on unseen categories and an open vocabulary of concepts without additional training. Extensive experiments demonstrate that OpenWorldSAM achieves state-of-the-art performance in open-vocabulary semantic, instance, and panoptic segmentation across multiple benchmarks, including ADE20k, PASCAL, ScanNet, and SUN-RGBD.
@@ 1 Introduction
    We introduce OpenWorldSAM, a unified interface that supports various open-vocabulary segmentation tasks. We propose an efficient language adapter with tie-breaker and cross-attention soft prompting, improving multi-object localization.
@@ 3 Methodology
Architecture overview. Figure 4 depicts our framework which comprises: (i) a hierarchical SAM2 image encoder that extracts image features, (ii) a multi-modal vision-language encoder that jointly ingests the image and text prompt, (iii) a lightweight MLP projector, (iv) learnable positional tie-breakers for multi-instance queries, (v) a soft prompting Transformer block that aligns text–image features with SAM2’s image features, and (vi) the SAM2 mask decoder producing final masks. Only a small language adapter with components (iii–v) is trained; all other backbones remain frozen.
@@
Multi-modal encoder. We leverage BEiT-3 [44] to encode the input description into a semantic embedding. Given an image $I$ and a text prompt $T$ (e.g., a category name or a referring expression), we feed both modalities into BEiT-3’s encoder to obtain joint visual–text embeddings. Concretely, tokens of $T$ and patch embeddings of a downsampled $I$ are concatenated and processed by BEiT-3, yielding a set of feature vectors $\{\mathbf{f}_{\text{[CLS]}},\mathbf{f}_{1},\dots\}$. We take the classification token $\mathbf{f}_{\text{[CLS]}}$ as a compact summary denoted as $\mathbf{p}_{\mathrm{lang}}$ of the prompt conditioned on the image content.
We adopt BEiT-3 because its early-fusion training on image-text pairs equips it with rich, bidirectional semantics—crucial for reasoning about unseen classes. Compared with CLIP-style contrastive image-text matching using only the features from the last encoder layers, BEiT-3 exposes finer cross-modal interactions. By embedding the text while it sees the image, the encoder already localizes the concept loosely (e.g., “giraffe” vs. “rock” in Figure 4) before any downstream segmentation, preventing the mask decoder from learning semantics from scratch.
Prompt projection. BEiT-3 emits 1,024-D tokens, whereas SAM’s prompt channels are 256-D. A two-layer MLP acts as a projector that (i) preserves the coarse semantics of $\mathbf{p}_{\mathrm{lang}}\in\mathbb{R}^{d_{1024}}$ and (ii) learns to highlight dimensions that are most useful for mask prediction: $\mathbf{u}=\mathrm{MLP}(\mathbf{p}_{\mathrm{lang}})\in\mathbb{R}^{d_{256}}.$
Positional tie-breaker and multi-instance queries generation. The projected visual-text embedding $\mathbf{u}$ captures what to segment but lacks awareness of how many instances exist and where they are. To enable multi-instance segmentation, we propose $K$ learnable positional tie-breaker vectors $\{\mathbf{t}_{1},\dots,\mathbf{t}_{K}\}\subset\mathbb{R}^{d_{256}}$ that perturb $\mathbf{u}$ into $K$ distinct queries:
@@
These perturbations serve two purposes: 1) Positional disambiguation: Each $\mathbf{t}_{i}$ nudges the query towards different spatial regions (Figure 3b), mimicking how human annotators might click different points to segment each zebra. 2) Instance diversity: The tie-breakers are optimized during training to maximize coverage of distinct instances, preventing query collapse. Conceptually these queries play the role of the “object queries” in DETR [45]. Crucially, they impose segmentation distinction for the same language semantics, making positional tie-breaking a novel and key feature for OpenWorldSAM. In practice $K=20$ covers $>$99% images in COCO [46]; for larger scenes $K$ can be increased trivially.
Soft-prompting via cross-attention. The perturbed queries $\{\mathbf{q}_{i}\}$ interact with SAM2’s image features through a 3-layer Transformer [47] in Figure 4c, which alternates self-attention (queries talk to each other, promoting diversity) and cross-attention (queries look at image features). SAM2’s image encoder follows a hierarchical vision Transformer (“Hiera” [48, 49]) that outputs three features $\{\mathbf{F}_{256\times 256},\mathbf{F}_{128\times 128},\mathbf{F}_{64\times 64}\}$ with $256^{2}$, $128^{2}$, and $64^{2}$ spatial resolutions, respectively. We operate on the level-3 features with $64^{2}$ resolution as they optimally balance precision for retainaing boundary details and computational efficiency (16×16$\times$16 × cheaper than full-resolution attention). They are also used by SAM2 for mask decoding by default [18]. The soft prompting Transformer computes $\mathbf{q}^{\prime}_{i}=\mathrm{CrossAttn}(\mathbf{q}_{i},\ \mathbf{F}_{64\times 64}),\;i=1,\dots,K,$ whose key/value inputs are the flattened level-3 features $\mathbf{F}_{64\times 64}\in\mathbb{R}^{4096\times 256}$. This step grounds the language-aware queries in SAM2’s high-resolution visual features, resolving ambiguities (e.g., distinguishing adjacent zebras by stripe patterns).
@@
Training. All heavy visual (Hiera) and vision-language (BEiT-3) encoders are kept frozen to preserve their pre-trained knowledge and avoid costly retraining. Only the MLP projector, tie-breakers, and the soft prompting Transformer are learnable. For each training sample and prompt, we match the $K$ predicted masks to the ground-truth masks of class $T$ via Hungarian matching [45], then apply a focal loss, encouraging precise segmentation of all instances described by the prompt. The tie-breakers $\mathbf{t}_{i}\in\mathbb{R}^{d_{256}}$ are implemented as learnable parameters randomly initialized from a normal distribution. During training, the Hungarian matching loss naturally encourages each $\mathbf{t}_{i}$ to specialize in different spatial regions. Notably, this mechanism requires no explicit supervision about instance counts.
@@
Optionally, we perform a two-stage inference. In this setup, masks obtained from the first inference stage are used as visual prompts fed back into SAM2’s mask decoder, which refines mask contours. Qualitative results in Appendix B demonstrate that this two-stage inference notably improves the precision of mask boundaries. However, quantitative analysis (Table 1) reveals that the second inference stage provides minimal improvements in segmentation metrics, suggesting it mainly enhances visual quality rather than overall accuracy.
@@ 4 Experiments
Implementation. We implement our model in PyTorch. We initialize the model with the public weights of SAM2-Hiera-Large and BEIT-3-Large. It is trained for 25 epochs on COCO-Stuff using the AdamW optimizer with a learning rate of 1e-4, batch size 8, on a single NVIDIA A100 GPU. Image resolution is set to 1024 for SAM2 and 224 for BEiT-3. Number of postional tie-breaks is set to 20 for COCO dataset. Our implementation details can be found in Appendix A.
@@ 4.2 Open-Vocabulary Segmentation Performance Analysis
Table 1: Zero-shot performance of open-vocabulary segmentation models across multiple benchmarks. For COCO, different methods use different supervisions of mask (m), class label (cls) and caption (cap). “ITP” indicates whether model uses image-text pairs/referring data. “DET” indicates extra detection data (e.g., Objects365, LVIS, OpenImages, etc.) “*” denotes the model has the capability for the task but does not have number reported. “-” means the model does not have the ability for the specific task. Purple color means a fully supervised approach, and tan a semi-supervised learning approach. Two-stage inference means we refine mask contours by re-prompting SAM using the raw mask predictions.
@@ 4.3 Referring Expression Segmentation Performance Analysis
Table 3: Referring segmentation performance (cIoU) comparison on RefCOCOg benchmark validation set between our proposed OpenWorldSAM and prior SOTA methods. We abbreviate the datasets: C (COCO), RC (RefCOCO/+), RCg (RefCOCOg), PL (PACO-LVIS), O365 (Objects365), V (Video segmentation datasets), OID (OpenImages Detection), VG (Visual Genome), ADE (ADE20K), PP (PASCAL-Part), PC (PASCAL-VOC). We compare model trainable parameters, model capabilities (OV seg (open-vocabulary segmentation) and Inter Seg (interactive segmentation)), and training data required. “*” denotes an estimate of the trainable parameters, since these models use LoRA [57] with rank-8/16 adapters for finetuning.
@@ 4.4 Ablation Studies
Optimal Training Strategy. In Table 4.3, we varied the trainable modules in OpenWorldSAM (thus varying total new parameters from  1.7M to  770M). We found in E5 that freezing BEiT-3 and training only the language adapter module (tie-breaker + cross-attention, 4.5M parameters) yields optimal performance (60.4 mIoU ADE-150). Notably, comparing E6 vs E5 and E4 vs E3, we found fine-tuning the entire BEiT-3 encoder (673M parameters) significantly degrades accuracy (mIoU drops from 60.4 to 23.6), likely due to underfitting on sparse category label prompts compared to its original web-scale pretraining.
@@ 5 Conclusion
Appendix A Experimental Settings
A.1 Pre-training
We implement our model in PyTorch, building on the Detectron2 [63] framework. We initialize the model with the public weights of SAM2-Hiera-Large [^1] https://github.com/facebookresearch/sam2 and BEIT-3-Large [^2] https://github.com/microsoft/unilm/tree/master/beit3. The model is pre-trained for 25 epochs on the COCO-2017 training split (104K images) [46], excluding the RefCOCOg-UMD validation subset [23]. We use the panoptic annotations, which provide pixel-accurate masks and category labels for all 132 thing and stuff classes.
Training is conducted on a single NVIDIA A100 (80 GB) GPU with a batch size of 8. Optimization employs AdamW (learning rate 1e-4). A step decay scheduler drops the learning rate by a factor of 0.1 at 89% and 96% of the total iterations. Compared with recent generalist models, our recipe is markedly more data-efficient (see Table 6).
Method Train Data (grouped by annotation types) Image Consumption
X-Decoder [15] COCO, RefC Cap4M 32, 1024 200M (50 Ep $\times$ 4M Img)
OpenSeeD [36] COCO, O365 32, 64 48M (30 Ep $\times$ 1.8M Img)
OpenWorldSAM COCO 8 2.50M (25 Ep $\times$ 0.104M Img)
A.2 Zero-Shot Evaluation
We evaluate semantic, instance, and panoptic segmentation in a zero-shot setting. For instance segmentation and panoptic segmentation, we apply confidence-score filtering to remove masks with scores below 0.7, followed by non-maximum suppression (NMS) with IoU threshold 0.5 to remove duplicate detections and retain distinct object instances. The confidence scores, originally termed “estimated IoU scores” in SAM [17, 18], are direct outputs from SAM2’s mask decoder. These scores were optimized during SAM2’s pre-training to select high-quality (i.e., confident) mask outputs.
Table 7: Open-Vocabulary Segmentation Benchmark Statistics.
Evaluation Dataset Scene type Annotations # Images # Classes
Semantic Instance Panoptic
ADE-150 common 2000 150
ADE-847 common 2000 847
Pascal Voc common 1449 20
Pascal Context-59 common 5105 59
Pascal Context-459 common 5105 459
SUN RGB-D in-door 5050 37
ScanNet-20 in-door 5436 20
ScanNet-40 in-door 5436 40
The open-vocabulary benchmark comprises 5 datasets covering 8 different segmentation tasks; statistics are summarized in Table 7. We show a comprehensive evaluation protocol for open-vocabulary segmentation in various vocabulary sizes and image domains.
A.3 Finetuning
For referring-expression segmentation we fine-tune the pre-trained checkpoint for 15 epochs. Because images from RefCOCOg were seen during pre-training (with category labels substituted for referring expressions ground truth), we adopt a conservative learning rate of 1e-5. We use a batch size of 8 during training.
Appendix B Qualitative Comparison on Two-Stage-Inference
During inference, we perform an optional two-stage inference. First, the model predicts multi-instance masks. These masks are then fed back as visual prompts, and SAM2’s mask decoder is run a second time to refine the contours. Figure 7 illustrates the visual improvement. However, quantitative gains are marginal across segmentation metrics (see Sec. 4.2 of the main paper), suggesting it mainly enhances visual quality rather than overall accuracy.
Appendix C Additional Zero-Shot Qualitative Results
(a) ADE20K-150
(b) ADE20K-857
(c) PASCAL Context-459
Appendix D Limitations - Outdoor Generalization
Model Evaluation Cityscapes BDD10K
mIoU AP PQ mIoU PQ
X-Decoder (L) [15] zero-shot 52.0 24.9 38.1 47.2 17.8
OpenWorldSAM zero-shot 39.4 10.1 26.4 31.3 15.6
OpenWorldSAM Finetune on Cityscapes 57.4 12.0 36.1 38.0 17.4
Hypotheses.
  1. (i)
  2. (ii)
    Resolution bottleneck. Cityscapes frames are $2048{\times}1024$. Rescaling to $1024{\times}1024$ (SAM default) reduces poles and traffic lights to nearly one pixel at the feature stride of 16×16$\times$16 ×. X-Decoder keeps an FPN branch at 8×8$\times$8 ×, preserving thin structures.
Take-away.
Appendix E Model Structure Details
Method Visual Backbone Base Model Text Encoder Image Size
Short Long
MSeg (B) [54] HRNet-W48 (65 M) HRNet-Seg 1024 1024
GroupViT (S) [27] ViT-S/16 (22 M) GroupViT Transformer 224 224
LSeg+ (B) [24] CLIP ViT-B/16 (86 M) DenseCLIP CLIP 512 512
ZegFormer (B) [55] CLIP ViT-B/16 (86 M) ZegFormer CLIP 640 640
OpenSeg (B) [11] ResNet-101 (45 M) OpenSeg CLIP/ALIGN 640 640
OVSeg (B) [12] CLIP ViT-B/16 (86 M) MaskFormer CLIP 640 640
MaskCLIP (L) [28] CLIP ViT-L/14 (307 M) MaskCLIP CLIP 1024 1024
X-Decoder [15] DaViT-L (196 M) X-Decoder CLIP 1024 1024
OpenSeeD [36] Swin-L (197 M) MaskDINO UniCL 1024 1024
SEEM [35] DaViT-L (196 M) X-Decoder CLIP 800 1333
APE (B) [16] ViT-L (307 M) DETA CLIP 1024 1024
PolyFormer (L) [58] Swin-L (197 M) PolyFormer BERT 1024 1024
UNINEXT (H) [32] ViT-H (632 M) DINO BERT 320$\sim$800 1333
PixelLM [60] CLIP ViT-L/14 (307 M) PixelLM LLaMA2-13B 448 448
LISA [19] SAM ViT-H (2.4 B) SAM Vicuna-7B 1024 1024
GLaMM [20] SAM ViT-H (2.4 B) SAM Vicuna-7B 1024 1024
u-LLaVA [21] SAM ViT-H (2.4 B) SAM Vicuna-7B 1024 1024
EVF-SAM [43] SAM ViT-H (2.4 B) SAM BEiT-3 1024 1024
EVF-SAM2 [43] SAM2 Hiera-L (224 M) SAM2 BEiT-3 1024 1024
OpenWorldSAM SAM2 Hiera-L (224 M) SAM2 BEiT-3 1024 1024
E.1 Possible Text Encoder Alternatives
We argue that the key ingredients for open-vocabulary segmentation are backbone-agnostic: any strong interactive segmenter can supply high-resolution mask decoding, while any pretrained vision-language encoder can provide semantics. What is missing is a lightweight adaptor that (i) aligns the two embedding spaces, (ii) scales to multiple object instances from a single text query, and (iii) preserves the efficiency that makes interactive segmentation attractive in the first place.
Our OpenWorldSAM is a general plug-in architecture that satisfies these desiderata while keeping all heavy backbones frozen. Although we instantiate the framework with SAM2 and BEiT-3 in this paper, neither component is required by design; alternative interactive decoders or vision-language encoders can be swapped in with only minor re-training of the adapter.
Early-fusion Transformers (drop-in closest to BEiT-3)
Dual-encoder Contrastive (text vector not image-conditioned)
Encoder–Decoder w/ Contrastive Head (pooled vector from decoder)
Query-former Hybrids (multiple vectors)
BLIP-2-OPT-2.7B [74] ViT + Q-Former + LLM 32$\times$256 1.1B native multi-query pool/average queries or extend SAM prompt len. 31.5

### v2

@@ Abstract
The ability to segment objects based on open-ended language prompts remains a critical challenge, requiring models to ground textual semantics into precise spatial masks while handling diverse and unseen categories. We present OpenWorldSAM, a framework that extends the prompt-driven Segment Anything Model v2 (SAM2) to open-vocabulary scenarios by integrating multi-modal embeddings extracted from a lightweight vision-language model (VLM). Our approach is guided by four key principles: i) Unified prompting: OpenWorldSAM supports a diverse range of prompts, including category-level and sentence-level language descriptions, providing a flexible interface for various segmentation tasks. ii) Efficiency: By freezing the pre-trained components of SAM2 and the VLM, we train only 4.5 million parameters on the COCO-stuff dataset, achieving remarkable resource efficiency. iii) Instance Awareness: We enhance the model’s spatial understanding through positional tie-breaker embeddings and cross-attention layers, enabling effective segmentation of multiple instances. iv) Generalization: OpenWorldSAM exhibits strong zero-shot capabilities, generalizing well on unseen categories and an open vocabulary of concepts without additional training. Extensive experiments demonstrate that OpenWorldSAM achieves state-of-the-art performance in open-vocabulary semantic, instance, and panoptic segmentation across multiple benchmarks. Code is available at https://github.com/GinnyXiao/OpenWorldSAM.
@@ 1 Introduction
    We introduce OpenWorldSAM, a unified interface that supports various open-vocabulary segmentation tasks. We propose an efficient language adapter with tie‑breaker and cross‑attention soft prompting, improving multi-object localization.
@@ 3 Methodology
Architecture overview. Figure 4 depicts our framework which comprises: (i) a hierarchical SAM2 image encoder that extracts image features, (ii) a multi-modal vision‐language encoder that jointly ingests the image and text prompt, (iii) a lightweight MLP projector, (iv) learnable positional tie‐breakers for multi‐instance queries, (v) a soft prompting Transformer block that aligns text–image features with SAM2’s image features, and (vi) the SAM2 mask decoder producing final masks. Only a small language adapter with components (iii–v) is trained; all other backbones remain frozen.
@@
Multi-modal encoder. We leverage BEiT-3 [44] to encode the input description into a semantic embedding. Given an image $I$ and a text prompt $T$ (e.g., a category name or a referring expression), we feed both modalities into BEiT‐3’s encoder to obtain joint visual–text embeddings. Concretely, tokens of $T$ and patch embeddings of a downsampled $I$ are concatenated and processed by BEiT‐3, yielding a set of feature vectors $\{\mathbf{f}_{\text{[CLS]}},\mathbf{f}_{1},\dots\}$. We take the classification token $\mathbf{f}_{\text{[CLS]}}$ as a compact summary denoted as $\mathbf{p}_{\mathrm{lang}}$ of the prompt conditioned on the image content.
We adopt BEiT‑3 because its early‑fusion training on image‑text pairs equips it with rich, bidirectional semantics—crucial for reasoning about unseen classes. Compared with CLIP‑style contrastive image-text matching using only the features from the last encoder layers, BEiT‑3 exposes finer cross‑modal interactions. By embedding the text while it sees the image, the encoder already localizes the concept loosely (e.g., “giraffe” vs. “rock” in Figure 4) before any downstream segmentation, preventing the mask decoder from learning semantics from scratch.
Prompt projection. BEiT‑3 emits 1,024‑D tokens, whereas SAM’s prompt channels are 256‑D. A two‑layer MLP acts as a projector that (i) preserves the coarse semantics of $\mathbf{p}_{\mathrm{lang}}\in\mathbb{R}^{d_{1024}}$ and (ii) learns to highlight dimensions that are most useful for mask prediction: $\mathbf{u}=\mathrm{MLP}(\mathbf{p}_{\mathrm{lang}})\in\mathbb{R}^{d_{256}}.$
Positional tie‑breaker and multi-instance queries generation. The projected visual-text embedding $\mathbf{u}$ captures what to segment but lacks awareness of how many instances exist and where they are. To enable multi-instance segmentation, we propose $K$ learnable positional tie-breaker vectors $\{\mathbf{t}_{1},\dots,\mathbf{t}_{K}\}\subset\mathbb{R}^{d_{256}}$ that perturb $\mathbf{u}$ into $K$ distinct queries:
@@
These perturbations serve two purposes: 1) Positional disambiguation: Each $\mathbf{t}_{i}$ nudges the query towards different spatial regions (Figure 3b), mimicking how human annotators might click different points to segment each zebra. 2) Instance diversity: The tie-breakers are optimized during training to maximize coverage of distinct instances, preventing query collapse. Conceptually these queries play the role of the “object queries” in DETR [45]. Crucially, they impose segmentation distinction for the same language semantics, making positional tie-breaking a novel and key feature for OpenWorldSAM. In practice $K=20$ covers >>99% images in COCO [46]; for larger scenes $K$ can be increased trivially.
Soft‐prompting via cross‐attention. The perturbed queries $\{\mathbf{q}_{i}\}$ interact with SAM2’s image features through a 3-layer Transformer [47] in Figure 4c, which alternates self‑attention (queries talk to each other, promoting diversity) and cross‑attention (queries look at image features). Each language‑aware query is refined on‑the‑fly by cross‑attention with the frozen SAM2 features. SAM2’s image encoder follows a hierarchical vision Transformer (“Hiera” [48, 49]) that outputs three features $\{\mathbf{F}_{256\times 256},\mathbf{F}_{128\times 128},\mathbf{F}_{64\times 64}\}$ with $256^{2}$, $128^{2}$, and $64^{2}$ spatial resolutions, respectively. We operate on the level-3 features with $64^{2}$ resolution as they optimally balance precision for retainaing boundary details and computational efficiency ($16\times$ cheaper than full‑resolution attention). They are also used by SAM2 for mask decoding by default [18]. The soft prompting Transformer computes $\mathbf{q}^{\prime}_{i}=\mathrm{CrossAttn}(\mathbf{q}_{i},\ \mathbf{F}_{64\times 64}),\;i=1,\dots,K,$ whose key/value inputs are the flattened level-3 features $\mathbf{F}_{64\times 64}\in\mathbb{R}^{4096\times 256}$. This step grounds the language-aware queries in SAM2’s high-resolution visual features, resolving ambiguities (e.g., distinguishing adjacent zebras by stripe patterns).
@@
Training. All heavy visual (Hiera) and vision‑language (BEiT‑3) encoders are kept frozen to preserve their pre‑trained knowledge and avoid costly retraining. Only the MLP projector, tie‑breakers, and the soft prompting Transformer are learnable. For each training sample and prompt, we match the $K$ predicted masks to the ground‐truth masks of class $T$ via Hungarian matching [45], then apply a focal loss, encouraging precise segmentation of all instances described by the prompt. The tie-breakers $\mathbf{t}_{i}\in\mathbb{R}^{d_{256}}$ are implemented as learnable parameters randomly initialized from a normal distribution. During training, the Hungarian matching loss naturally encourages each $\mathbf{t}_{i}$ to specialize in different spatial regions. Notably, this mechanism requires no explicit supervision about instance counts.
@@
Optionally, we perform a two-stage inference. In this setup, masks obtained from the first inference stage are used as visual prompts fed back into SAM2’s mask decoder, which refines mask contours. Qualitatively, two-stage inference improves the precision of mask boundaries for correct predictions (Appendix B). However, quantitative analysis (Table 1) reveals that the second inference stage provides minimal improvements in segmentation metrics, suggesting it mainly enhances visual quality rather than overall accuracy.
@@ 4 Experiments
Implementation. We implement our model in PyTorch. We initialize the model with the public weights of SAM2-Hiera-Large and BEIT-3-Large. It is trained for 25 epochs on COCO-Stuff using the AdamW optimizer with a learning rate of 1e‑4, batch size 8, on a single NVIDIA A100 GPU. Image resolution is set to 1024 for SAM2 and 224 for BEiT-3. Number of postional tie-breaks is set to 20 for COCO dataset. Our implementation details can be found in Appendix A.
@@ 4.2 Open-Vocabulary Segmentation Performance Analysis
Table 1: Zero-shot performance of open-vocabulary segmentation models across multiple benchmarks. For COCO, different methods use different supervisions of mask (m), class label (cls) and caption (cap). “ITP” indicates whether model uses image-text pairs/referring data. “DET” indicates extra detection data (e.g., Objects365, LVIS, OpenImages, etc.) “*” denotes the model has the capability for the task but does not have number reported. “-” means the model does not have the ability for the specific task. Purple color means a fully supervised approach, and tan a semi-supervised learning approach. Two-stage inference means we refine mask contours by re-prompting SAM using the raw mask predictions. Bold entries indicate the best performance.
@@ 4.3 Referring Expression Segmentation Performance Analysis
Table 3: Referring segmentation performance (cIoU) comparison on RefCOCOg benchmark validation set between our proposed OpenWorldSAM and prior SOTA methods. We abbreviate the datasets: C (COCO), RC (RefCOCO/+), RCg (RefCOCOg), PL (PACO-LVIS), O365 (Objects365), V (Video segmentation datasets), OID (OpenImages Detection), VG (Visual Genome), ADE (ADE20K), $P$ (PASCAL-Part), PC (PASCAL-VOC). We compare model trainable parameters, model capabilities (OV seg (open-vocabulary segmentation) and Inter Seg (interactive segmentation)), and training data required. “*” denotes an estimate of the trainable parameters, since these models use LoRA [58] with rank-8/16 adapters for finetuning.
@@ 4.4 Ablation Studies
Optimal Training Strategy. In Table 4.3, we varied the trainable modules in OpenWorldSAM (thus varying total new parameters from  1.2M to  770M). We found in E5 that freezing BEiT-3 and training only the language adapter module (tie-breaker + cross-attention, 4.5M parameters) yields optimal performance (60.4 mIoU ADE-150). Notably, comparing E6 vs E5 and E4 vs E3, we found fine-tuning the entire BEiT-3 encoder (673M parameters) significantly degrades accuracy (mIoU drops from 60.4 to 23.6), likely due to underfitting on sparse category label prompts compared to its original web-scale pretraining.
@@ 5 Conclusion
Acknowledgement. This work was supported in part by CoCoSys, a JUMP2.0 center sponsored by DARPA and SRC, the National Science Foundation (CAREER Award, Grant #2312366, Grant #2318152), the DARPA Young Faculty Award and the DoE MMICC center SEA-CROGS (Award #DE-SC0023198).